Montserrat Volcano Observatory, Montserrat, West Indies

Scientific Report 78
23 November 1997

This report covers the 14 day period from 00:00 on 9 November to 00:00 (local time) on 23 November, 1997.


Activity has been at a relatively low level over the past two weeks. Dome observations have been scarce due to poor weather conditions. The Galway's lobe appears to be continuing as the focus of activity but extrusion rates may have slowed. Rockfall signals and very small amplitude hybrid earthquakes have dominated the seismic records.

Visual Observations

Visual observations of the dome have been severely limited during this period due to poor weather. An observation flight on 11th November revealed that the southern lobe had grown quite significantly almost doubling in size over 3 days. The Galway's lobe of the October 22nd dome has rapidly filled the collapse scar produced by the collapses on the 4th and 6th of November but the extrusion rate of the lava may have slowed since the October dome first appeared. Ash and steam venting has still been occurring from the cleft or vent between the two lobes. The ash clouds produced from the venting commonly reach 6,000ft a.s.l. and have mostly been drifting westwards with ashfall over Plymouth. Rockfalls spalling off the new growth area continue to occur predominantly in the Galway's area with the northern lobe remaining mostly inactive. The Galway's rockfalls have been eroding chutes down the southern side of the dome and accumulating a thick mass of deposits in the Galway's Soufriere region.


For the first part of the reporting period seismicity remained at a high level with large numbers of hybrid earthquakes and episodes of tremor. There was a small pyroclastic flow on the 9th November but apart from this the hybrids did not lead up to any external activity on the volcano. The hybrids were smaller in amplitude than those recorded earlier in November and continued to become smaller over the first few days of the reporting period. From the 14th onwards activity was dominated by rockfall signals although there continued to be large numbers of very low amplitude hybrids - many of which did not trigger the networks. These hybrids were scattered throughout the day and not grouped in swarms.

Table 1: Earthquake types

These earthquake counts are of events that triggered the broadband network's event recording system between 00:00 and 00:00 each day (local time).

Date		VT	Hybrid	LP	Dome RF		LPRF*	HYRF*

09 Nov 97	0	646	10	12		0	0
10 Nov 97	0	330	34	23		0	0
11 Nov 97	0	51	4	41		1	4
12 Nov 97	1	375	19	51		9	3
13 Nov 97	0	132	8	49		3	0
14 Nov 97	0	54	4	65		0	0
15 Nov 97	2	12	2	75		0	0
16 Nov 97	9	24	15	87		0	0
17 Nov 97	0	19	11	56		1	0
18 Nov 97	0	3	4	111		1	0
19 Nov 97	0	23	26	79		6	0
20 Nov 97	0	122	21	87		4	1
21 Nov 97	0	216	27	111		10	0
22 Nov 97	1	35	16	76		2	1

* LPRF: LP earthquake followed by rockfall signal. HYRF: Hybrid earthquake followed by rockfall signal. The LPs, hybrids and rockfalls in these signals are also counted in their respective columns.

Table 2: Swarms

Start    	Stop	     	# of vts	# of hybrids	# of lps
09/11/97 03:12	11/11/97 03:01	    0		996		44
12/11/97 03:23	13/11/97 05:53	    0		412		9

Ground Deformation

No surveys were undertaken during this period.

Volume Measurements

A survey was carried out on the 17th November of the fan and deposits in the lower reaches of the White River valley. The total deposits in the White River valley has a volumes of 13.6 million cubic metres which is an increased of 5.5 million since the last survey on 15th May. Most of these deposits are considered to have been emplaced during the two collapses on the 4th and 6th of November. This total does not include the volume of deposits that have accumulated recently in the upper reaches of the valley - this has remained unsurveyed due to high ash levels. The relative sizes of the collapses is less well constrained but best estimates suggest that of the 4th was ~1.8 million and that of the 6th was ~3.4 million.

Environmental Monitoring

Dust Trak sampling has been carried out at four fixed sites. Each value is an average of the PM10 concentration over approximately 24 hrs. The values at the fixed sites have been low over this period except for the values at the Catholic school which sometimes recorded raised levels. This is due to the large amount of human activity at this site and the fact that it is near a main road. Towards the end of this reporting period the averages recorded at the three sites not including the school all had remarkably similar average PM10 concentrations each day.

A new Dust Trak site was established at Mango Drive in Woodlands on 16 November to replace the Runaway site.

Table 3 : PM10 (concentration in mg/m3)

Site		09 Nov 	10 Nov 	11 Nov 	12 Nov 	13 Nov 	14 Nov 	15 Nov 	
Runaway		0.034	0.018	0.017	0.015	0.012			
St. Peter's	0.031	0.018	0.012	0.014	0.013		0.025	
CPS		0.039	0.054	0.041	0.065	0.074		0.065	
MVO north	0.026	0.013	0.014	0.01	0.013		0.026	

Site		16 Nov 	17 Nov 	18 Nov 	19 Nov 	20 Nov 	21 Nov 	22 Nov 	
Mango		0.014	0.012	0.011	0.013	0.012	0.016	0.035	
St. Peter's	0.017	0.014	0.015	0.012	0.011	0.014	0.035	
CPS		0.023	0.046	0.031	0.045	0.027	0.021	0.06	
MVO north	0.013	0.01	0.011	0.012	0.012	0.016	0.034	

These values are approximate 24 hr averages of the PM10 concentration.

24 Hour averages:
<0.05 mg m-3	Low
0.05-0.1 mg m-3	Raised
0.1-0.3 mg m-3	Very High
>0.3 mg m-3	Alert

Runaway:        The Dust Trak is outside on the patio of a villa in Olveston.
St. Peter's:    The Dust Trak is outside on the balcony of a villa in St Peter's,
			called St. Peter's Place.
CPS:            Catholic Primary School The Dust Trak is outside in the area under the
			roof where some of the children have classes. The school is
			in Palm Loop near Woodlands.
MVO north:      The new Volcano observatory on Mongo Hill near St.John's, the Dust trak is
			on the third story of the building with the sampling tube stuck
			out of the window.
Mango: 		The Dust Trak is outside on the verandah of a villa in Mango Drive in Woodlands

MVO Staff Changes:
Dr Willy Aspinall arrived to take over as chief scientist
Dr Sayadul Arafin, Mr Chan Ramsingh and Mr Godfrey Almorales (Seismic Research Unit in Trinidad)
Dr Gill Norton from the BGS arrived to take over as deputy chief scientist.

Chief scientist Dr Keith Rowley finished his tour of duty and has returned to Trinidad
Deputy chief scientist Dr Sue Loughlin (from BGS)

HRH Duke of York (on 22nd November)

Montserrat Volcano Observatory